OpenStudy (anonymous):

Hello, I know the answer is D. Can someone tell me the procedure used to get the answer? Find the indicated probability. Round to three decimal places. A test consists of 10 true/false questions. To pass the test a student must answer at least 6 questions correctly. If a student guesses on each question, what is the probability that the student will pass the test? A. 0.828 B. 0.172 C. 0.205 D. 0.377

jimthompson5910 (jim_thompson5910):

do you have a calculator like a TI-83 or TI-84?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

TI-84 Plus

OpenStudy (amistre64):

A test consists of 10 true/false questions. To pass the test a student must answer at least 6 questions correctly. If a student guesses on each question, what is the probability that the student will pass the test? P(x=6) + P(x=7) + P(x=8) + P(x=9) + P(x=10) is at least 6 right

OpenStudy (amistre64):

open your distribution menu; 2nd vars i beleive, and pick binomCDF

OpenStudy (anonymous):

Okay, I have opened binomcdf

OpenStudy (amistre64):

or since P(t) = .5 and P(f) = .5 the binomial thrm get us: \[\sum_{n=6}^{10}\binom{10}{n}(.5)^{n}(.5)^{10-n}\] \[\sum_{n=6}^{10}\binom{10}{n}(.5)^{10}\] \[(.5)^{10}\sum_{n=6}^{10}\binom{10}{n}\] should work as well

OpenStudy (amistre64):

binomCDF(n,p,x) are the inputs n=10, p=.5, x=5 why? binom counts up from x=0 to x=k we want from (k+1) to 10 soo 1 - binom(10,.5, 5) gets us there